Circular No. 6987

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

1998 ML_{14}

M. Hicks and P. Weissman, Jet Propulsion Laboratory, write: “We report BVRI photometry of 1998 ML_{14} (MPEC 1998-M26, 1998-P17) taken with the Las Campanas 2.54-m telescope of the Carnegie Institution of Washington, over three nights (July 12–14 UT). After subtraction of solar colors, we determined relative magnitudes of B − R = +0.203 ± 0.005, V − R = +0.088 ± 0.004, I − R = +0.084 ± 0.004, consistent with S-type asteroids. Our R-band photometry was best phased together into a doublepeaked light curve with an amplitude of 0.12 mag, a synodic period of 14.98 ± 0.06 hr and a mean absolute Kron-Cousins R magnitude of 16.93 ± 0.01, assuming a value for the slope parameter of G = 0.25.”

SATELLITES OF NEPTUNE C. Roddier, F. Roddier, J. E. Graves, O. Guyon, and M. J. North-

cott, Institute for Astronomy, University of Hawaii (UH), report: “Infrared (1.72-μm) images of Neptune were taken on July 6 with the UH adaptive optics system (Hokupa’a) mounted on the Canada-France-Hawaii telescope. Neptune VIII (Proteus), VII (Larissa), and VI (Galatea) were detected and observed from July 6.490 to 6.552 UT. Whereas Proteus and Larissa were found to be at their expected position, Galatea was found to be 5^{o} ± 1^{o} ahead of the predicted position (8.6 min early), a difference possibly due to its interaction with Neptune’s Adams ring (cf. Ferrari and Brahic 1994, Icarus 111, 193; Hanninen and Porco 1997, Icarus 126, 1).”

EV LACERTAE F. Favata, Astrophysics Division, European Space Agency/ESTEC;

and S. Sciortino, G. Micela, and A. Maggio, Palermo Astronomical Observatory, report: “We observed the dMe flare star EV Lac in x-rays with the ASCA satellite during July 13.213–15.069 UT. An unusually strong flaring event took place between July 13.833 and 13.948. A preliminary analysis of the ASCA data from the SIS and GIS detectors shows that the event is of unprecedented intensity for single-star flares, with the peak x-ray luminosity exceeding the optical V -band luminosity by a factor of ∼ 5 (i.e., reaching a value higher than the bolometric photospheric luminosity). Observers who happened to be monitoring EV Lac at any other wavelength (especially optical and radio) at the time of the x-ray observation are invited to contact the above authors (ffavata@astro.estec.esa.nl).”

1998 August 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green